OpenStudy (yacoub1993):

Find the standard form of the equation of the ellipse and give the location of its foci.

OpenStudy (yacoub1993):

@satellite73

OpenStudy (yacoub1993):

A. x^2/36 +y^2/49 = 1 foci at (- sqrt13 , 0) and ( sqrt13 , 0) B. x^2/49 - y^2/36 = 1 foci at (- sqrt13 , 0) and ( sqrt13 , 0) C. x^2/49 + y^2/36= 1 foci at (- sqrt13 , 0) and ( sqrt13 , 0) D. x^2/49 + y^2/36 = 1 foci at (-7, 0) and ( 7, 0)

OpenStudy (yacoub1993):

I choose D and it was marked wrong

OpenStudy (yacoub1993):

@thomaster Help me Please

OpenStudy (yacoub1993):

@zepdrix Help me Please

OpenStudy (yacoub1993):

@sweetburger Help me Please